John is 8 years older than his sister. In 3 years he will be twice as old as she is. How old are they now?
Let x represent the age of John's sister now

Answer by stanbon(75887)

John is 8 years older than his sister. In 3 years he will be twice as old as she is. How old are they now?
Let s represent the age of John's sister now
----
Equations:
j = s + 8
j+3 = 2(s+8)
----
Substitute for "j" and solve for "s":
s+8+3 = 2s+16
s+11 = 2s+16
s = -5 (Sis's age now)
Comment: That's the answer but that doesn't make any sense.
--
Cheers,
Stan H.